# Medformer: A Multi-Granularity Patching Transformer for Medical Time-Series Classification

### v1

@@ Appendix E Complexity Analysis
Appendix F Discussion
F.1 Comparision with Other Multi-Granularity Methods
MTST [45] and Pathformer [47] differ from our Medformer in three significant aspects: (1) Patching & Embedding MTST and Pathformer utilize single-channel patching, presupposing channel independence. In contrast, Medformer employs multi-channel patching to capture potential channel correlations. (2) Architecture Structure MTST and Pathformer adopt a parallel multi-branch architecture, where each branch comprises independent attention modules tailored to specific granularities. Conversely, Medformer utilizes a shared attention module across different granularities, thereby significantly reducing model complexity. (3) Granularity Interactions MTST assimilates multi-granularity information by concatenating outputs from different branches, while Pathformer uses adaptive pathways for weighted aggregation of these outputs without any inter-granularity interactions within the attention modules. In contrast, Medformer introduces a novel inter-granularity attention mechanism specifically designed for granularity interaction, thereby effectively integrating multi-granularity information.
@@ F.1 Comparision with Other Multi-Granularity Methods
F.2 Limitations and Future Works
The design of Medformer enables the input of various patch lengths, offering both benefits and challenges. The ability to choose variable patch lengths suggests the potential for discovering configurations that outperform those with uniform lengths, as evidenced by our experiments (See Appendix D.2 and Appendix C). Nonetheless, not all combinations of patch lengths yield optimal results; some configurations may perform worse than uniform patch lengths. Consequently, this flexibility necessitates meticulous tuning of patch lengths as hyperparameters. Future research could explore the development of mechanisms that automatically select the most effective patch lengths, optimizing for the most relevant granularities. Additionally, investigating the performance of large medical time series models across different datasets is also an interesting direction.
Appendix G Broader Impacts

### v2

@@ Appendix F Complexity Analysis
Appendix G Discussion
G.1 Comparision with Other Multi-Granularity Methods
MTST [53] and Pathformer [55] differ from our Medformer in three significant aspects: (1) Patching & Embedding MTST and Pathformer utilize single-channel patching, presupposing channel independence. In contrast, Medformer employs multi-channel patching to capture potential channel correlations. (2) Granularity Interactions MTST assimilates multi-granularity information by concatenating outputs from different branches, while Pathformer uses adaptive pathways for weighted aggregation of these outputs without any inter-granularity interactions within the attention modules. In contrast, Medformer introduces a novel inter-granularity attention mechanism specifically designed for granularity interaction, thereby effectively integrating multi-granularity information.
@@ G.1 Comparision with Other Multi-Granularity Methods
Appendix H Broader Impacts